# A Monocular Wide-FoV SLAM System Based on Point Uncertainty and Pose Uncertainty

### v1

@@ III-B Kannala-Brandt Fisheye Model
The format of the Kannala-Brandt fisheye model (KB model) [42] is highly similar to the Taylor camera model, except that the KB model drops all the even terms in the polynomial of Eq. (6) to get closer to the real physical model.
@@ IV Problem Statement
Many VO and visual SLAM methods use Bundle Adjustment (BA) to estimate the pose of sensor $\mathbf{T}_{i}$ at time $i$ and the position of the $j$th map point $\mathbf{p}_{j}$, where $i=1,...,M$ and $j=1,...,N$. These variables to estimate are represented as $x=\{\mathbf{T}_{0},...,\mathbf{T}_{M},\mathbf{p}_{1},...,\mathbf{p}_{N}\}$, where one subset $\{\mathbf{T}_{i},\mathbf{p}_{j}\}$ indicates a potential observation constraint between $i$th pose and $j$th map point. That constraint will be explicit when sensors receive measurement $\mathbf{u}_{ij}$ of $j$th map point at time $i$ and be expressed as
@@
The existing standard formulation of SLAM in practice derives from the essential paper of Lu and Milios [57], in which SLAM is summarized as a maximum likelihood estimation problem to obtain optimal variables to estimate, expressed as
@@
where the specific format of ${\rm P}(\mathbf{u}_{ij}|\mathbf{T}_{i},\mathbf{p}_{j})$ is determined by probability distribution of sensor’s measurement noise $\mathbf{n}_{ij}$ in Eq. (7), which is usually assumed as a zero-mean Gaussian distribution with covariance matrix ${\Sigma}_{ij}$. Then, Eq. (9) can be rewritten as
@@
$=\mathop{\arg\min}\limits_{\mathbf{T}_{i},\mathbf{p}_{j}}||\pi(\mathbf{T}_{i}\cdot\mathbf{p}_{j})-\mathbf{u}_{ij}||_{{\Sigma}_{ij}^{-1}}^{2},$
where the notion $||e||_{{\Sigma}^{-1}}^{2}=e^{T}{\Sigma}^{-1}e$. It is notable that the covariance matrix ${\Sigma}_{ij}$ is only relevant to the sensor’s actual measurement $\mathbf{u}_{ij}$ and irrelevant to the ideal observation calculation $\pi(\mathbf{T}_{i}\cdot\mathbf{p}_{j})$.
@@ IV-A Point Uncertainty
This subsection briefly introduces the role that 3D point uncertainty plays in motion estimation and explains how it influences estimation.
Eq. (10) works well in a situation where all the variables to estimate are involved in the BA process while existing SLAM and VO methods utilize the technique of local map optimization considering the accumulative quantity of variables along with time and restrictive computing resources. The application of a local map limits the number of variables involved in BA to ensure the efficiency of optimization. In the meantime, it introduces the transformation of partial variables’ properties, especially when the sensor is a fisheye or a PAL camera owning an ultra-wide overlap area, which past wide-FoV SLAM frameworks ignore. The much wider overlap area between frames increases the probability that the map points generated in the past are observed by the newest frame. Thus, the wide FoV results in a greater proportion of well-estimated past map points in the local map. In order to avoid computational costs and help estimate the pose of the latest frame, these past map points are treated as measurement results rather than estimation variables. This shift in data properties has become more common with the introduction of wide-FoV sensors. Therefore, the uncertainty of these past map points should be introduced and the conditional probability part in Eq. (9) also needs to be modified. Considering a observation constraint between a pose to estimate $\mathbf{T}_{i}$ and a past map point $\mathbf{p}_{p}$, we have
@@
where there are two measurements $\mathbf{u}_{ip}$ with random noise $\mathbf{n}_{ip}\sim N(0,{\Sigma}_{ip})$ and $\mathbf{p}_{p}\sim N(\hat{\mathbf{p}_{p}},{\Sigma}_{3\times 3,p})$. ${\Sigma}_{3\times 3,p}$ denotes the the uncertainty of past map point $\mathbf{p}_{p}$. Then, the covariance ${\Sigma}_{ip}^{\prime}$ of ${\rm P}(\mathbf{u}_{ip},\mathbf{p}_{p}|\mathbf{T}_{i})$ can be calculated according to Eq. (5) and Eq. (7):
${\Sigma}_{ip}^{\prime}=\nabla_{\hat{\mathbf{p}_{p}}}\pi\cdot{\Sigma}_{3\times 3,p}\cdot(\nabla_{\hat{\mathbf{p}_{p}}}\pi)^{T}+{\Sigma}_{ip},\$ (12)
@@
$\nabla_{\hat{\mathbf{p}_{e}}}\pi$ $=\frac{\partial\pi(\mathbf{T}_{i}\cdot\hat{\mathbf{p}_{p}})}{\partial\hat{\mathbf{p}_{p}}}$ (13)
@@
$\mathbf{T}_{i}^{*}$ $=\mathop{\arg\min}\limits_{\mathbf{T}_{i}}-{\rm log}({\rm P}(\mathbf{u}_{ip},\mathbf{p}_{p}|\mathbf{T}_{i})$ (14)
$=\mathop{\arg\min}\limits_{\mathbf{T}_{i}}||\pi(\mathbf{T}_{i}\cdot\hat{\mathbf{p}_{p}})-\mathbf{u}_{ip}||_{{\Sigma}_{ip}^{\prime-1}}^{2}.$
The point uncertainty ${\Sigma}_{3d,p}$ is determined by historical observation of map point $\mathbf{p}_{p}$. For one historical observation $\mathbf{u}_{kp}$ at time $k$, the 3D observation error $\mathbf{r}_{kp}$ of $\mathbf{p}_{p}$ is defined as
$\mathbf{r}_{kp}=\mathbf{R}_{k}^{T}\cdot(\mathbf{T}_{k}\cdot\mathbf{p}_{p}-\mathbf{T}_{k}\cdot\mathbf{p}_{p}\cdot{\pi}^{-1}(\mathbf{u}_{kp})),$ (15)
@@
where $K$ denotes a subset of all the timestamps of the map point $\mathbf{p}_{p}$ historical observation, and $N_{K}$ denotes the count of historical observation.
@@ IV-B Pose Uncertainty
$=\mathop{\arg\min}\limits_{\mathbf{p}_{j}}||\pi(\mathbf{T}_{p}\cdot\hat{\mathbf{p}_{j}})-\mathbf{u}_{pj}||_{{\Sigma}_{pj}^{\prime\prime-1}}^{2}.$
and the corresponding covariance ${\Sigma}_{pj}^{\prime\prime}$ is
${\Sigma}_{pj}^{\prime\prime}=\nabla_{\hat{\xi_{p}}}\pi\cdot{\Sigma}_{6\times 6,p}\cdot(\nabla_{\hat{\xi_{p}}}\pi)^{T}+{\Sigma}_{pj},\$ (18)
@@
which involves the exponential mapping of the Lie group and applies a perturbation model [58].
@@ V System Overview
A pipeline of our system is schematically given as Fig. 4. P2U-SLAM is established on the monocular mode of ORB-SLAM3 [23] and mainly includes initialization, tracking, local bundle adjustment, and loop closing. The construction, application, and maintenance of point and pose uncertainties will be integrated into the main framework of P2U-SLAM to provide robustness and positioning accuracy. And ORB features are selected to apply in feature matching of the whole above process.
@@ V-A Initialization
As stated in PALVO [20], the map points observed by the wide-FoV camera are barely possible to be coplanar. Thus essential matrix is utilized to resolve the process of structure from motion (SFM). The essential matrix satisfies the epipolar constraints.
@@
The eight-point algorithm with a RANSAC scheme [59] is applied to exclude outliers of keypoint matches. However, the epipolar in the non-perspective projection camera model as mentioned in Sec. III is not a straight line but a “curve”, which makes it difficult to calculate reprojection errors in polar geometry. Therefore, the reprojection error is redesigned as the reprojection of vertical distance between $\mathbf{\overline{p}}^{i2}$ and polar plane, which is used to be the inlier-checking metric of RANSAC. After successful initialization, the system will begin to estimate the pose using the constraint relationship between map points and the current frame.
@@ V-B Tracking with Point Uncertainty
The tracking module aims at real-time updating of the agent pose. In this module, as shown in Fig. 3, the only variable to estimate is the pose $\mathbf{T}_{i}$ of the new frame and all of the map points $\mathbf{p}_{j}$ with matched 2D-features $\mathbf{u}_{ij}$ are measurement results, so according to Eq. (14), we have
$\mathbf{T}_{i}^{*}=\mathop{\arg\min}\limits_{\mathbf{T}_{i}}\sum_{j\in E}||\pi(\mathbf{T}_{i}\cdot\hat{\mathbf{p}_{j}})-\mathbf{u}_{ij}||_{{\Sigma}_{ij}^{\prime-1}}^{2},$ (22)
where $E$ is the set of map points observed in the new frame, and point uncertainty influences the direction of optimization as information matrix ${\Sigma}_{ij}^{\prime-1}$ in Eq. (12). The tracking module is where real-time performance truly shines, and with point uncertainty, P2U-SLAM enhances the accuracy of pose estimation. This approach allows for better handling of past map points as measurements, leading to more precise tracking.
@@ V-C Local BA with Pose Uncertainty
The local BA module further optimizes the pose of local keyframes and all visible map points to provide reliable visual constraints for subsequent tracking. In the process of local BA, the whole map points are set variables to estimate while keyframes are divided into fixed keyframes (observation results) and local keyframes (variables to estimate) as shown in Fig. 3. Therefore, the optimization can be expressed as
$\mathbf{T}_{I_{1}}^{*},\mathbf{p}^{*}=\mathop{\arg\min}\limits_{\mathbf{T}_{I_{1}},\mathbf{p}}\sum_{j}($ $\sum_{i_{1}\in I_{1}}||\pi(\mathbf{T}_{i_{1}}\cdot\mathbf{p}_{j})-\mathbf{u}_{i_{1}j}||_{{\Sigma}_{i_{1}j}^{-1}}^{2}$ (23)
@@ V-D Loop Closing
In the thread of loop closing, all ORB features obtained from keyframes have been transformed to BoW vectors with DBoW2 [48]. The candidate loop keyframes are filtered by comparing BoW similarity with the current keyframe. False positives in loop detection tend to appear more frequently with a wide-FoV camera because of its much bigger overlapping area. Therefore, the epipolar constraints mentioned in Eq. (21) will also be utilized to check between the candidate loop keyframe and the current keyframe. Once a true positive loop closure is detected, the relative pose of the current keyframe from the candidate loop keyframe will be calculated and all uncertainty of poses and points will be updated. Finally, a global BA is processed to optimize keyframe poses and map points.
@@ VI Experiments
In this section, the proposed P2U-SLAM is evaluated on two public datasets, including the PALVIO dataset [26] and the TUM-VI benchmark [60]. The PALVIO dataset, recorded with two Panoramic Annular Lens cameras and an IMU sensor, is collected on an aerial vehicle as shown in Fig. 1, providing a comprehensive dataset for assessing SLAM in indoor environments. The TUM-VI benchmark, on the other hand, features fisheye inputs that mimic the viewpoint of a vehicle or drone in motion, offering a set of challenging sequences to test the robustness of SLAM systems in various dynamic conditions. Notably, P2U-SLAM is a monocular visual SLAM system, and all the compared methods are set to the monocular mode for a fair comparison except that IMU data input is imperative. The estimated trajectories of SLAM and VIO systems are aligned with ground truth using a Sim(3) transformation. All experiments have been run on an AMD Core r5-4650g CPU, with 16GB memory, using only the CPU for computing.
Seq. ID02
Seq. ID04
Seq. ID06
(a) Translation and Rotation Error
(b) Top Trajectory
Figure 5: Examples of error analyses and top trajectories of different SLAM and VIO systems on the PALVIO dataset [26]. The details of the a) Translation and Rotation Error subfigures illustrate that P2U-SLAM not only has a low average error but also exhibits a more concentrated and stable error distribution. In the b) Top Trajectory subfigures, the clear superiority of P2U-SLAM in overlap with the ground truth trajectory over VINS-Mono and LF-VIO demonstrates its advantage of low absolute trajectory error, while its higher trajectory smoothness compared to ORB-SLAM3 indicates the stability of its real-time tracking. This is attributed to the innovative integration of point- and pose uncertainty within the P2U-SLAM framework, which effectively mitigates the impact of weak feature correspondences and enhances the overall reliability of the system.
[…]
@@
VINS-Mono [9] RPEt 4.118 5.936 2.068 1.098 / 2.231 / 3.567 9.747 2.372 3.114*
@@
ORB-SLAM3 (mono) [23] RPEt / 2.132 1.604 1.935 1.861 2.009 2.010 2.239 2.662 / 1.645*
@@
    Average error of the successful sequences. Systems that did not complete all sequences are denoted by *.
@@
    The data of SVO2.0 are referenced from [26].
[…]

### v2

@@ III-B Kannala-Brandt Fisheye Model
The format of the Kannala-Brandt fisheye model (KB model) [kannala2006generic] is highly similar to the Taylor camera model, except that the KB model drops all the even terms in the polynomial of Eq. (6) to get closer to the real physical model.
@@ IV Problem Statement
Many VO and visual SLAM methods use Bundle Adjustment (BA) to estimate the pose of sensor $\mathbf{T}_{i}$ at time $i$ and the position of the $j$th map point $\mathbf{p}_{j}$, where i=1,,Mi=1,...,M and j=1,,Nj=1,...,N. These variables to estimate are represented as $x=\{\mathbf{T}_{0},...,\mathbf{T}_{M},\mathbf{p}_{1},...,\mathbf{p}_{N}\}$, where one subset $\{\mathbf{T}_{i},\mathbf{p}_{j}\}$ indicates a potential observation constraint between $i$th pose and $j$th map point. That constraint will be explicit when sensors receive measurement $\mathbf{u}_{ij}$ of $j$th map point at time $i$ and be expressed as
@@
The existing standard formulation of SLAM in practice derives from the essential paper of Lu and Milios [lu1997globally], in which SLAM is summarized as a maximum likelihood estimation problem to obtain optimal variables to estimate, expressed as
@@
where the specific format of ${\rm P}(\mathbf{u}_{ij}|\mathbf{T}_{i},\mathbf{p}_{j})$ is determined by the probability distribution of the sensor’s measurement noise $\mathbf{n}_{ij}$ in Eq. (7), which is usually assumed as a zero-mean Gaussian distribution with covariance matrix ${\Sigma}_{ij}$. Then, Eq. (9) can be rewritten as
@@
[…]
where the notion $e||_{{\Sigma}^{-1}}^{2}=e^{T}{\Sigma}^{-1}e$.
It should be noted that in Eq.(10), both $\mathbf{T}_{i}$ and $\mathbf{p}_{j}$ are the targets we aim to solve for, and only $\mathbf{u}_{ij}$ represents the observed results. The original intention of Eq.(10) is to find a set of $\mathbf{T}_{i}$ and $\mathbf{p}_{j}$ such that the likelihood probability is maximized. Therefore, the covariance ${\Sigma}_{ij}$ of the likelihood distribution is only related to the observed result $\mathbf{u}_{ij}$, and is independent of $\mathbf{T}_{i}$ and $\mathbf{p}_{j}$. Meanwhile, the result of Eq.(10) is exactly the optimization objective function for global Bundle Adjustment (BA) in the visual SLAM problem.
@@ IV-A Point Uncertainty
This subsection introduces the role that 3D point uncertainty plays in motion estimation and explains how it influences estimation.
Eq. (10) works well in a situation where all the variables to estimate are involved in the BA process while existing SLAM and VO methods utilize the technique of local map optimization considering the cumulative quantity of variables along with time and restrictive computing resources. The application of a local map limits the number of variables involved in BA to ensure the efficiency of optimization. In the meantime, it introduces the transformation of partial variables’ properties, especially when the sensor is a fisheye or a PAL camera owning an ultra-wide overlap area, which past wide-FoV SLAM frameworks ignore. The much wider overlap area between frames increases the probability that the map points generated in the past are observed by the newest frame. Thus, the wide FoV results in a greater proportion of well-estimated past map points in the local map. In order to avoid computational costs and help estimate the pose of the latest frame, these past map points are treated as measurement results rather than estimation variables. This shift in data properties has become more common with the introduction of wide-FoV sensors. Therefore, the uncertainty of these past map points should be introduced and the conditional probability part in Eq. (9) also needs to be modified. Considering a observation constraint between a pose to estimate $\mathbf{T}_{i}$ and a past map point $\mathbf{p}_{p}$, we have
@@
where the data property of $\mathbf{p}_{p}$ is converted from the variable to be estimated in Eq.(10) into observable data. There are two measurements. One is $\mathbf{u}_{ip}$ with random noise $\mathbf{n}_{ip}\sim N(0,{\Sigma}_{ip})$ and the other is $\mathbf{p}_{p}\sim N(\hat{\mathbf{p}_{p}},{\Sigma}_{3\times 3,p})$. ${\Sigma}_{3\times 3,p}$ denotes the the uncertainty of past map point $\mathbf{p}_{p}$. Then, the covariance ${\Sigma}_{ip}^{\prime}$ of ${\rm P}(\mathbf{u}_{ip},\mathbf{p}_{p}|\mathbf{T}_{i})$ can be calculated according to Eq. (5) and Eq. (7):
${\Sigma}_{ip}^{\prime}=\nabla_{\hat{\mathbf{p}_{p}}}\pi\cdot{\Sigma}_{3\times 3,p}\cdot(\nabla_{\hat{\mathbf{p}_{p}}}\pi)^{T}+{\Sigma}_{ip},\\$ (12)
@@
$\nabla_{\hat{\mathbf{p}_{p}}}\pi$ $=\frac{\partial\pi(\mathbf{T}_{i}\cdot\hat{\mathbf{p}_{p}})}{\partial\hat{\mathbf{p}_{p}}}$ (13)
@@
$\mathbf{T}_{i}^{*}$ $=\mathop{\arg\min}\limits_{\mathbf{T}_{i}}-{\rm log}({\rm P}(\mathbf{u}_{ip},\mathbf{p}_{p}|\mathbf{T}_{i}))$ (14)
[…]
The point uncertainty ${\Sigma}_{3\times 3,p}$ is determined by historical observation of map point $\mathbf{p}_{p}$. For one historical observation $\mathbf{u}_{kp}$ at time $k$, the 3D observation error $\mathbf{r}_{kp}$ of $\mathbf{p}_{p}$ is defined as
[…]
@@
where $K$ denotes a subset of all the timestamps of the map point $\mathbf{p}_{p}$ historical observation, and $N_{K}$ denotes the count of historical observations.
@@ IV-B Pose Uncertainty
[…]
where the data property of $\mathbf{T}_{p}$ is converted from the variable to be estimated in Eq.(10) into observable data. The distribution of the pose measurement values is represented in the form of Lie algebra, which is $\xi_{p}\sim N(\hat{\xi_{p}},{\Sigma}_{6\times 6,p})$. ${\Sigma}_{6\times 6,p}$ denotes the the uncertainty of past pose’s Lie algebra $\xi_{p}$ . And the corresponding covariance ${\Sigma}_{pj}^{\prime\prime}$ is
${\Sigma}_{pj}^{\prime\prime}=\nabla_{\hat{\xi_{p}}}\pi\cdot{\Sigma}_{6\times 6,p}\cdot(\nabla_{\hat{\xi_{p}}}\pi)^{T}+{\Sigma}_{pj},\\$ (18)
@@
which applies a perturbation model [1972Introduction] and involves the exponential mapping of the Lie group, whose specific form is as follows,
@@
Similarly to Eq. (15), the pose uncertainty ${\Sigma}_{6\times 6,p}$ is determined by the mapping of the two-dimensional historical observations in the Lie algebra space of the pose. Unlike the calculation of 3D map point observation error, where we have a camera back-projection function that can directly obtain 3D observation results from 2D observations to simplify the calculation, in the acquisition of pose observation error, there is no such convenient function for converting between 2D point observations and 6D pose observations. Therefore, here we directly use the pseudo-inverse of the gradient matrix of the 2D projection with respect to the pose Lie algebra and the 2D observation error to calculate the pose observation error. For one historical observation $\mathbf{u}_{ph}$ at time $p$, the 6D observation error $\mathbf{r}_{ph}$ of $\mathbf{\xi}_{p}$ is defined as
$\mathbf{r}_{ph}=(\nabla_{\hat{\xi_{p}}}\pi(\mathbf{T}_{p}\cdot\mathbf{p}_{h}))^{-1}\cdot(\mathbf{u}_{ph}-{\pi}(\mathbf{T}_{p}\cdot\mathbf{p}_{h})),$ (21)
where $(\nabla_{\hat{\xi_{p}}}\pi)^{-1}$ denotes the pseudo-inverse of $\nabla_{\hat{\xi_{p}}}\pi$
$(\nabla_{\hat{\xi_{p}}}\pi)^{-1}=(\nabla_{\hat{\xi_{p}}}\pi)^{T}\cdot(\nabla_{\hat{\xi_{p}}}\pi\cdot(\nabla_{\hat{\xi_{p}}}\pi)^{T})^{-1},$ (22)
and the pose uncertainty ${\Sigma}_{6\times 6,p}$ is
${\Sigma}_{6\times 6,p}=\frac{1}{N_{H}-1}\sum_{h\in H}\mathbf{r}_{ph}\cdot{\mathbf{r}_{ph}}^{T},$ (23)
where $H$ denotes a subset of all the historical observations $\{\mathbf{u}_{ph}|h\in H\}$ in keyframe at time $p$ whose pose is $\mathbf{T}_{p}$, and $N_{H}$ denotes the count of historical observations.
@@ V System Overview
A pipeline of our system is schematically given as Fig. 3. P2U-SLAM is established on the monocular mode of ORB-SLAM3 [campos2021orb] and mainly includes initialization, tracking, local bundle adjustment, and loop closing. The construction, application, and maintenance of point and pose uncertainties have been integrated into the main framework of P2U-SLAM to provide robustness and positioning accuracy. And ORB features are selected to apply in the feature matching of the whole above process.
@@ V-A Initialization
As stated in PALVO [chen2019palvo], the map points observed by the wide-FoV camera are barely possible to be coplanar. Thus essential matrix is utilized to resolve the process of structure from motion (SFM). The essential matrix satisfies the epipolar constraints.
@@
The eight-point algorithm with a RANSAC scheme [longuet1981computer] is applied to exclude outliers of keypoint matches. However, the epipolar in the non-perspective projection camera model as mentioned in Sec. III is not a straight line but a “curve”, which makes it difficult to calculate reprojection errors in polar geometry. Therefore, the reprojection error is redesigned as the reprojection of vertical distance between $\mathbf{\overline{p}}^{i2}$ and the polar plane, which is used to be the inlier-checking metric of RANSAC. After successful initialization, the system will begin to estimate the pose using the constraint relationship between map points and the current frame.
@@ V-B Tracking with Point Uncertainty
The tracking module aims at real-time updating of the agent’s pose. In this module, as shown in Fig. 4, the only variable to estimate is the pose $\mathbf{T}_{i}$ of the new frame and all of the map points $\mathbf{p}_{j}$ with matched 2D-features $\mathbf{u}_{ij}$ are measurement results, so according to Eq. (14), we have
[…]
where $E$ is the set of map points observed in the new frame, and point uncertainty influences the direction of optimization by acting as the information matrix ${\Sigma}_{ij}^{\prime-1}$ in Eq. (12). The tracking module is where real-time performance truly shines, and with point uncertainty, P2U-SLAM enhances the accuracy of pose estimation. This approach allows for better handling of past map points as measurements, leading to more precise tracking.
Seq. ID02
Seq. ID04
Seq. ID08
(a) Translation and Rotation Error
(b) Top Trajectory
Figure 5: Examples of error analyses and top trajectories of different SLAM and VIO systems on the PALVIO dataset [wang2022lf]. The details of the a) Translation and Rotation Error subfigures illustrate that P2U-SLAM not only has a low average error but also exhibits a more concentrated and stable error distribution. In the b) Top Trajectory subfigures, the clear superiority of P2U-SLAM in overlap with the ground truth trajectory over VINS-Mono and LF-VIO demonstrates its advantage of low absolute trajectory error, while its higher trajectory smoothness compared to ORB-SLAM3 indicates the stability of its real-time tracking. This is attributed to the innovative integration of point- and pose uncertainty within the P2U-SLAM framework, which effectively mitigates the impact of weak feature correspondences and enhances the overall reliability of the system.
@@ V-C Local BA with Pose Uncertainty
The local BA module further optimizes the pose of local keyframes and all visible map points to provide reliable visual constraints for subsequent tracking. This means that the pose of local keyframes and all visible map points are our optimization targets, and their uncertainties are exactly what we aim to reduce. Therefore, those uncertainties cannot be used as the covariance of observed data to participate in the optimization. That is to say, only the uncertainty of the pose of the fixed keyframe can be used for Local BA. In the process of local BA, the whole map points are set variables to estimate, while keyframes are divided into fixed keyframes (observation results) and local keyframes (variables to estimate) as shown in Fig. 4. Therefore, the optimization can be expressed as
[…]
@@ V-D Loop Closing
In the thread of loop closing, all ORB features obtained from keyframes have been transformed to BoW vectors with DBoW2 [galvez2012bags]. The candidate loop keyframes are filtered by comparing BoW similarity with the current keyframe. False positives in loop detection tend to appear more frequently with a wide-FoV camera because of its much bigger overlapping area. Therefore, the epipolar constraints mentioned in Eq. (24) will also be utilized to check between the candidate loop keyframe and the current keyframe. Once a true positive loop closure is detected, the relative pose of the current keyframe from the candidate loop keyframe will be calculated and all uncertainty of poses and points will be updated. Finally, a global BA is processed to optimize keyframe poses and map points.
@@ VI Experiments
In this section, the proposed P2U-SLAM is evaluated on two public datasets, including the PALVIO dataset [wang2022lf] and the TUM-VI benchmark [schubert2018tum]. The PALVIO dataset, recorded with two Panoramic Annular Lens cameras and an IMU sensor, is collected on an aerial vehicle as shown in Fig. 1, providing a comprehensive dataset for assessing SLAM in indoor environments. The TUM-VI benchmark, on the other hand, features fisheye inputs that mimic the viewpoint of a vehicle or drone in motion, offering a set of challenging sequences to test the robustness of SLAM systems in various dynamic conditions. Notably, P2U-SLAM is a monocular visual SLAM system, and all the compared methods are set to the monocular mode for a fair comparison except that IMU data input is imperative. The estimated trajectories of SLAM and VIO systems are aligned with the ground truth using a Sim(3) transformation. All experiments have been run on an AMD Core r5-4650g CPU, with 16GB memory, using only the CPU for computing.
TABLE I: Performance comparison on the public PALVIO dataset [wang2022lf] (ATE in $m$, RPEt in $\%$, RPEr in degree/mdegree/m)1.
@@
VINS-Mono [qin2018vins] RPEt 4.118 5.936 2.068 1.098 / 2.231 / 3.567 9.747 2.372 3.114*
@@
LF-VISLAM [wang2023lfvislam] RPEt 6.869 3.026 2.621 1.707 1.448 1.687 1.479 1.835 1.516 3.493 2.568
RPEr 0.398 0.740 0.199 0.111 0.139 0.171 0.145 0.348 0.362 0.300 0.291
ATE 0.360 0.398 0.460 0.276 0.181 0.102 0.263 0.190 0.171 0.339 0.274
ORB-SLAM3 (mono) [campos2021orb] RPEt / 2.132 1.604 1.935 1.861 2.009 2.010 2.239 2.662 / 1.645*
@@
    Average error of the successful sequences. Systems that did not complete all sequences are marked with *.
@@
    The data of SVO2.0 are referenced from [wang2022lf].
Figure 6: The robustness of several algorithms on the PALVIO dataset is demonstrated, mainly through three metrics: a) RMS ATE (m), b) RMS RPEt ($\%$), and c) RMS RPEr (degree/m). Each algorithm runs a total of ten times on each sequence, with the results of each run represented by the color squares. A higher red component in the square indicates a worse result in the corresponding metric test, while a greater blue component signifies better performance. An incomplete run is represented by a blank white square. It should be noted that the incomplete runs of ORB-SLAM3 (mono) are slightly different from those of VINS-MONO. In the monocular mode, ORB-SLAM3 has no IMU input. Once the visual constraints fail, the loss of positioning cannot be recovered, and there is no difference in the results of repeated experiments on the same failure sequence. However, VINS-MONO still has the motion constraints of the IMU when the visual constraints fail. Therefore, there will be inconsistent results in repeated experiments on the same sequence.